### v2

@@
Abstract
Pareto Set Learning (PSL) is popular as an efficient approach to obtaining the complete optimal solution in Multi-objective Learning (MOL). A set of optimal solutions approximates the Pareto set, and its mapping is a set of dense points in the Pareto front in objective space. However, some current methods face a challenge: how to make the Pareto solution is diverse while maximizing the hypervolume value. In this paper, we propose a novel method to address this challenge, which employs Stein Variational Gradient Descent (SVGD) to approximate the entire Pareto set. SVGD pushes a set of particles towards the Pareto set by applying a form of functional gradient descent, which helps to converge and diversify optimal solutions. Additionally, we employ diverse gradient direction strategies to thoroughly investigate a unified framework for SVGD in multi-objective optimization and adapt this framework with an annealing schedule to promote stability. We introduce our method, SVH-MOL, and validate its effectiveness through extensive experiments on multi-objective problems and multi-task learning, demonstrating its superior performance.
keywords:
@@ Abstract
1 Introduction
@@
Multi-objective optimization (MOO) is a critical area of research in the field of optimization, focusing on problems that involve multiple conflicting objectives. Unlike traditional single-objective optimization, where the goal is to find a single optimal solution, MOO seeks to find a set of solutions that represent the trade-offs between the different objectives. MOO has been successfully applied across diverse fields, from energy system design [1] to healthcare treatment planning [2], demonstrating its versatility in balancing conflicting objectives. In machine learning, multi-objective problems play a significant role in various applications, including recommender systems [3, 4, 5], where they help balance multiple conflicting criteria, and multi-task learning [6, 7], where they optimize performance across multiple related tasks simultaneously. Furthermore, addressing multi-objective problems in real-world scenarios can be computationally demanding [8], particularly in high-dimensional settings [9], as it involves assessing multiple conflicting objectives, which increases computational costs.
@@
A common method for addressing multi-objective problems is to transform them into a single-objective problem using scalarization. This method assigns weights to each objective, representing their trade-offs and relative importance. The optimization process then finds a single solution corresponding to a specific weight vector. Thus, for each weight vector, we obtain only a single solution within a specific subregion of the objective space, which is defined by the chosen weighting scheme and the trade-offs between objectives. This subregion represents a portion of the Pareto front that aligns with the given weight vector. As a result, the disadvantage is how choose the desired weight. MOEA/D [10] (Multi-Objective Evolutionary Algorithm based on Decomposition) is a widely used multi-objective optimization algorithm that decomposes a multi-objective problem into several single-objective subproblems and optimizes them simultaneously. MOEA/D enhances solution diversity by decomposing the problem into multiple subproblems, allowing for a more distributed exploration of the solution space.
@@
The current survey [11] has classified MOO algorithms for multi-task learning into three approaches following the number of optimal solutions:
@@
  • Finding a single Pareto solution: A single optimal solution is obtained, representing a solution with the contribution of each objective appropriately, ensuring that the final solution is not overly biased toward any particular objective. This balance is typically achieved through techniques such as loss balancing [12, 13, 14] and gradient-based adjustments [6, 15, 16], which dynamically regulate the influence of each objective during the optimization process.
@@
  • Finding a finite set of Pareto solutions: This approach provides users with a diverse set of options to choose from based on their specific requirements. Generally, preference vector-based methods are used to divide the problem into subproblems, where each solution is found by optimizing a specific subproblem defined by a preference vector. Furthermore, methods such as Pareto Multi-Task Learning (PMTL) [17] leverage this strategy by evenly distributing preference vectors across the objective space to capture different trade-off regions. Alternatively, algorithms like Exact Pareto Optimal (EPO) [18] Search use techniques such as gradient descent with controlled ascent to precisely navigate the Pareto front and locate solutions that exactly satisfy the user-specified preferences.
@@
  • Finding an infinite set of Pareto solutions: This approach seeks to compute the complete Pareto front—that is, a continuum of Pareto optimal solutions that are continuously and densely distributed in the objective space. The hypernetwork-based method [19, 20] obtains a good Pareto front by mapping the preference vector space directly to the optimal solution space. Moreover, this hypernetwork-based framework significantly reduces the need for training multiple separate models for various preference settings. Instead, a single model can be employed to generate a diverse set of solutions, making it especially suitable for complex multi-task learning scenarios where balancing objectives is critical.
@@
In our research, we investigate the third approach discussed in the context of Pareto set learning methods, which involves utilizing hypernetwork-based strategies. We observe that the current method faces a major challenge, which is that the Pareto set is difficult to diversify. In PHN-HVI [20], the authors leverage hypervolume maximization [21] to derive the optimal solution while incorporating a penalty function to promote diversity among the resulting Pareto solutions. Using Hypervolume maximization as a gradient descent direction helps to find the optimal solution, but similar to MGDA, it tends to move towards a single optimal point. Therefore, in PHN-HVI, the incorporation of a penalty function promotes a more diverse spread of solutions along the Pareto front. However, this diversification can come at the cost of convergence precision, potentially steering the optimization process away from the most optimal points in the objective space.
@@
To tackle this challenge, we propose a novel approach for efficiently learning the complete Pareto set while capturing its inherent diversity. Our method is inspired by the previous study MOO-SVGD [22], which employs Stein Variational Gradient Descent (SVGD) [23] to approximate the Pareto set. However, MOO-SVGD results in a discrete set of solutions. In contrast, our approach is designed to learn a continuous representation of the Pareto front, ensuring that all possible trade-offs are effectively captured. Our framework, SVH-MOL, has the following contributions:
  • Firstly, we introduce an innovative framework that integrates hypernetwork-based techniques with Stein Variational Gradient Descent for multi-objective learning.
  • Secondly, we investigate several variations of the gradient descent algorithm within the SVGD framework to fully capture and optimize the capabilities of our approach. Additionally, we employ an annealing schedule to encourage diversity in the Pareto set.
  • Thirdly, we evaluated our framework across a broad spectrum of applications, ranging from traditional multi-objective optimization (MOO) tasks to large-scale multi-task learning problems. In our experiments, we demonstrated that our approach not only effectively captures the intricate trade-offs inherent in MOO scenarios but also scales robustly to complex, real-world multi-task settings.
2 Related Work
Multi-objective Optimisation (MOO)
@@ 2 Related Work
MOO is a fundamental challenge in machine learning, where it aims to find a set of optimal solutions that each solution represents the best trade-off between multiple conflicting objectives. Evolutionary algorithms are popular to find a set of diverse Pareto optimal solutions in a single run, such as VEGA [24], NSGA-III [25]. This approach often faces challenges in high-dimensional spaces and require significant computational resources. In contrast, gradient-based methods have emerged as more efficient alternatives, particularly in differentiable settings.This method leverages the gradients of the objective functions to guide the search toward Pareto-optimal solutions. Scalarization techniques, such as the weighted sum method [26] and the Tchebyshev method [27], provide an efficient approach for transforming multiple objectives into a single aggregated objective, enabling the use of gradient-based optimization methods.
@@
PSL emerges as a proficient means of approximating the entire Pareto front; it employs a hypernetwork [28] to learn a mapping between a preference vector and an optimal solution in a single training. Many studies proposed different gradient descent methods to optimize the combined-objective function, such as PHN-LS and PHN-EPO [19]. [20] use maximizing the hypervolume indicator to guide the search for the optimal region. One framework for controllable Pareto front learning employs the completed scalarization functions proposed by [29]. PSL has been effectively applied to the optimization of multiple black-box functions [8], where it can efficiently suggest the next experiment to conduct in a laboratory setting.
@@
SVGD is a gradient method for approximating a target distribution, which was originally proposed for Bayesian inference [23]. SVGD iteratively transports a set of particles to approximate a target distribution by minimizing the Kullback-Leibler (KL) divergence. However, [30] obverses that particles of SVGD tend to collapse to modes of the target distribution, and this particle degeneracy phenomenon becomes more severe with higher dimensions. This issue was explained in this study [31], the authors compared the SVGD update with the gradient descent on the maximum mean discrepancy (MMD). To address this challenge, [32] employs an annealing schedule to solve the inability of the particles to escape from local modes and the inefficiency in reproducing the density of the different regions. Selecting an appropriate kernel remains a key challenge, as the performance of the algorithm is highly dependent on the choice of kernel [33]. Multiple-kernel learning (MKL) enables the combination of multiple kernels into a single, unified kernel, providing greater flexibility in the learning process [34]. In MOO, the first study MOO-SVGD [35] adapted SVGD to profile the Pareto front while maintaining diversity, providing an efficient alternative to traditional methods.
3 Preliminaries
3.1 Multi-objective optimization
@@
$\min_{x\in\mathcal{X}}\mathcal{F}(x)=\begin{bmatrix}f_{1}(x)\\ f_{2}(x)\\ \vdots\\ f_{m}(x)\end{bmatrix},$ (1)
@@
3.2 Pareto Set Learning
Figure 1: Pareto Set Learning Framework
Pareto Set Learning (PSL) offers an efficient approach to obtaining the entire Pareto set in a single training phase, contrasting with traditional methods that require varying hyperparameters to approximate only a few Pareto-optimal points. PSL employs a hypernetwork [28] to generate a diverse set of candidate solutions. During training, the hypernetwork’s parameters are optimized by directly minimizing the multi-objective optimization (MOO) objective. Consequently, the trained hypernetwork produces a dense set of solutions that closely approximate the true Pareto front.
@@ 3.2 Pareto Set Learning
$\begin{split}\phi^{*}=\arg\min_{\phi}\mathbb{E}_{r\sim\text{Dir}(\alpha)}\mathbf{s}\left(\mathcal{F}\left(\mathbf{x}_{r}\right),r\right)\\ \text{s.t. }\mathbf{x}_{r}=\mathbf{h}(r,\phi^{*})\in X_{E},\,\mathbf{h}\left(\mathcal{P},\phi^{*}\right)\equiv X_{E}\end{split}$ (2)
where Hypernetwork $h:\mathcal{P}\times\mathbb{R}^{q}\to\mathbb{R}^{n}$, $q$ is the number of parameters of hypernetwork, and Dir$(\alpha)$ denoting the Dirichlet distribution and $r$ being the preference vector. During the optimization process, we update the hypernetwork parameter $\phi$ as follows:
$\phi_{t+1}=\phi_{t}-\xi\nabla_{\phi}\mathbf{s}\left(\mathcal{F}\left(\mathbf{x}_{r}\right),r\right),$ (3)
@@
$\phi^{*}=\arg\min_{\phi}\mathbb{E}_{r\sim Dir(\alpha)}s(\mathcal{L}(\cdot,\theta_{r}),r)$ (4)
@@
3.3 Stein Variational Gradient Descent (SVGD)
[23] proposed a particle-based inference algorithm, SVGD, to approximate the target distribution by applying the gradient information in the updating process. SVGD starts at a set of initial points $\{x_{i}^{0}\}_{i=1}^{n}\in\mathbb{R}^{d}$ drawn by a prior distribution $p(x)$, which is pushed to the target points $\{x^{*}_{i}\}_{i=1}^{n}$ belonging to the target distribution $q(x)$, following the iterative:
$\begin{split}x_{i}^{t+1}\leftarrow x_{i}^{t}+\epsilon\phi^{*}(x_{i}^{t}),\forall i=1,...,n,\\ \phi^{*}_{k}=\arg\max_{\phi\ \in\mathcal{B}_{k}}\left\{-\frac{d}{d\epsilon}\mathbf{K}\mathbf{L}(q_{[\epsilon\phi]}||p)\bigg{|}_{\epsilon=0}\right\}\end{split}$ (5)

### v3

@@
Abstract.
Pareto Set Learning (PSL) is an efficient approach for approximating the complete set of optimal solutions in Multi-objective Learning (MOL). By learning a set of solutions that map to a dense Pareto front in objective space, PSL aims to recover the underlying Pareto set. However, most existing methods focus heavily on convergence toward optimality, often neglecting solution diversity. To explicitly balance convergence and diversity, we propose Stein Variational Hypernetwork for MOL (SVH-MOL), a novel framework that incorporates Stein variational updates into Pareto set learning. SVH-MOL updates solutions via two complementary components: (i) a driving term that guides particles toward the Pareto set, and (ii) a repulsive term that encourages diversity among solutions. These two terms inherently compete, making stable and effective learning challenging. To address this issue, we introduce an annealing schedule that adaptively controls the relative strength of each term during training. Extensive experiments on synthetic multi-objective benchmarks and real-world multi-task learning problems demonstrate that SVH-MOL achieves a better trade-off between convergence and diversity than existing methods.
@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: JDSjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8
1. Introduction
@@
Multi-objective optimization (MOO) is a critical area of research in the field of optimization, focusing on problems that involve multiple conflicting objectives. Unlike traditional single-objective optimization, where the goal is to find a optimal solution(s) for a single objective function, MOO seeks to find a set of solutions that represent the trade-offs between the different objectives.MOO has been adopted across many domains, from energy system design (Marler2004) to healthcare treatment planning (Craft2006), demonstrating its versatility in balancing conflicting objectives. In machine learning, multi-objective problems play a significant role in various applications, including recommender systems (milojkovic2019multi; jannach2022multi; zaizi2023multi), where they help balance multiple conflicting criteria, and multi-task learning (sener2018multi; crawshaw2020multi), where they optimize performance across multiple related tasks simultaneously. Furthermore, addressing multi-objective problems in real-world scenarios can be computationally demanding (lin2022pareto), particularly in high-dimensional settings (nguyen2024high), as it involves assessing multiple conflicting objectives, which increases computational costs. Recently, training large language models (LLMs) has integrated MOO to efficiently transfer downstream tasks, which are related to multi-task fine-tuning (liu2024mftcoder) and multi-objective alignment (wu2023fine; zhou2024beyond).
@@
A common method for addressing multi-objective problems is to transform them into a single-objective problem using scalarization. This method assigns weights to each objective, representing their trade-offs and relative importance. The optimization process then finds a single solution corresponding to a specific weight vector. As a result, the disadvantage is how to choose the desired weight. MOEA/D (4358754) (Multi-Objective Evolutionary Algorithm based on Decomposition) is a widely used multi-objective optimization algorithm that decomposes a multi-objective problem into several single-objective subproblems and optimizes them simultaneously. MOEA/D enhances solution diversity by decomposing the problem into multiple subproblems, allowing for a more distributed exploration of the solution space.
@@
A current survey (chen2025gradient) has classified MOO algorithms for multi-task learning into three approaches following the number of optimal solutions:
@@
  • Finding a single Pareto solution: A single optimal solution is obtained, representing a solution with the contribution of each objective appropriately, ensuring that the final solution is not overly biased toward any particular objective. This balance is typically achieved through techniques such as loss balancing (liu2019end; kendall2018multi; lin2024smooth) and gradient-based adjustments (sener2018multi; chen2018gradnorm; yu2020gradient), which dynamically regulate the influence of each objective during the optimization process.
@@
  • Finding a finite set of Pareto solutions: This approach provides users with a diverse set of options to choose from based on their specific requirements. Generally, preference vector-based methods are used to divide the problem into subproblems, where each solution is found by optimizing a specific subproblem defined by a preference vector. Furthermore, methods such as Pareto Multi-Task Learning (PMTL) (lin2019pareto) leverage this strategy by evenly distributing preference vectors across the objective space to capture different trade-off regions. Alternatively, algorithms like Exact Pareto Optimal (EPO) (mahapatra2020multi) Search use techniques such as gradient descent with controlled ascent to precisely navigate the Pareto front and locate solutions that exactly satisfy the user-specified preferences.
@@
  • Finding an infinite set of Pareto solutions: This approach seeks to compute the complete Pareto front—that is, a continuum of Pareto optimal solutions that are continuously and densely distributed in the objective space. The hypernetwork-based method (navon2020learning; Hoang2023) obtains a good Pareto front by mapping the preference vector space directly to the optimal solution space. Moreover, this hypernetwork-based framework significantly reduces the need for training multiple separate models for various preference settings. Instead, a single model can be employed to generate a diverse set of solutions, making it especially suitable for complex multi-task learning scenarios where balancing objectives is critical.
@@
In the pioneer work, MOO-SVGD (liu2021profiling), Stein Variational Gradient Update (SVGD) is integrated to find a finite Pareto set. However, this adaptation suffers from several key weaknesses. The gradient descent direction is not in control, as it follows the Multiple Gradient Descent Algorithm (MGDA). It might be harder for MGDA to reach the global optimal region than other methods. Many weak Pareto points are obtained, which are not the desirable trade-off in MOO. Furthermore, MOO-SVGD in multi-task learning is not feasible. In fact, a large number of samples in training hinders the parameter optimization efficiency because of generating many architectures for all samples together. In other study, PHN-HVI (Hoang2023), the authors leverage hypervolume maximization (deist2021multiobjectivelearningpredictpareto) to derive the optimal solution while incorporating a penalty function to promote diversity among the resulting Pareto solutions. Using Hypervolume maximization as a gradient descent direction helps to find the optimal solution, but similar to MGDA, it tends to move towards a single optimal point. Therefore, in PHN-HVI, the incorporation of a penalty function promotes a more diverse spread of solutions along the Pareto front. However, this diversification can come at the cost of convergence precision, potentially steering the optimization process away from the most optimal points in the objective space.
To tackle the above issues, in our research, we investigate the third approach discussed in the context of Pareto set learning methods, which involves utilizing hypernetwork-based strategies. We leverage a hypernetwork to generate samples efficiently. We can adapt many strategies for controlling the Pareto front in training hypernet; thus, our approach is able to find a global Pareto set, which guarantees both convergence and diversity. For the hypernetwork-based approach, the issue about parameter optimization is mitigated, as the hypernetwork outputs the architecture for each sample, respectively.
@@
We propose a novel approach for efficiently learning the complete Pareto set while capturing its inherent diversity, which employs Stein Variational Gradient Descent (SVGD) (liu2016stein) to approximate the Pareto set. Our approach is designed to learn a continuous representation of the Pareto front, ensuring that all possible trade-offs are effectively captured. While the integration of Stein Variational Gradient Descent with hypernetworks was initalliy explored in our prior work SVH-PSL (Nguyen_Dinh_Nguyen_Hoang_Le_2025) for expensive multi-objective optimization, this paper introduces several fundamental advancements that specifically address challenges in multi-objective learning:
  • Novel Annealing Framework for MOL: We propose a specialized annealing schedule that dynamically balances convergence and diversity throughout training, unlike the static optimization in SVH-PSL designed for expensive black-box functions.
  • Controllable Scalarization Strategies: We systematically investigate three distinct scalarization methods (Linear, Tchebyshev, Smooth Tchebyshev) within the SVGD framework and analyze their intricate interactions with the repulsive force term—an aspect unexplored in SVH-PSL.
  • Large-Scale Multi-Task Learning Focus: While SVH-PSL targeted low-dimensional expensive optimization, SVH-MOL specifically addresses the scalability challenges in high-dimensional multi-task learning with deep neural networks, where parameter efficiency and training stability are paramount.
  • Comprehensive Theoretical and Empirical Analysis: We provide extensive ablation studies on hyperparameter sensitivity, computational complexity, and convergence behavior, along with discussions on limitations and future directions—significantly expanding beyond the scope of SVH-PSL.
2. Related Work
Multi-objective Optimization (MOO)
@@ 2. Related Work
MOO is a fundamental challenge in machine learning, where it aims to find a set of optimal solutions that each solution represents the best trade-off between multiple conflicting objectives. Evolutionary algorithms are popular to find a set of diverse Pareto optimal solutions in a single run, such as VEGA (schaffer2014multiple), NSGA-III (deb2013evolutionary). This approach often faces challenges in high-dimensional spaces and require significant computational resources. In contrast, gradient-based methods have emerged as more efficient alternatives, particularly in differentiable settings. This method leverages the gradients of the objective functions to guide the search toward Pareto-optimal solutions. Scalarization techniques, such as the weighted sum method (YANG2014197) and the Tchebyshev method (steuer1983interactive), provide an efficient approach for transforming multiple objectives into a single aggregated objective, enabling the use of gradient-based optimization methods.
@@
PSL emerges as a proficient means of approximating the entire Pareto front; it employs a hypernetwork (ha2016hypernetworks) to learn a mapping between a preference vector and an optimal solution in a single training. Many studies proposed different gradient descent methods to optimize the combined-objective function, such as PHN-LS and PHN-EPO (navon2020learning). PHN-HVI (Hoang2023) use maximizing the hypervolume indicator to guide the search for the optimal region. One framework for controllable Pareto front learning employs the completed scalarization functions proposed by (tuan2024framework). To exploit the geometric characteristics of the Pareto set in PSL, PSL-HV (zhang2023hypervolume) is introduced, which derives Pareto solutions that closely align with the polar angle under mild conditions. In a recent study, PO-PSL (haishan2024preference) focuses on the optimization of the preference vector instead of employing randomly generated vectors. This approach initiates a bilevel optimization problem in which it operates a preference-optimized term first. PO-PSL can lead to faster convergence to the true Pareto front. PSL has been effectively applied to optimization of multiple black-box functions (lin2022pareto), where it can efficiently suggest the next experiment to perform in a laboratory setting. The first work for Multi-Objective Reinforcement Learning is called PSL-MORL (liu2025pareto), which apply Pareto set learning to output a personalized policy network for each preference.
@@
SVGD (liu2016stein) is a gradient method for approximating a target distribution, which was originally proposed for Bayesian inference. SVGD iteratively transports a set of particles to approximate a target distribution by minimizing the Kullback-Leibler (KL) divergence. However, as observed in (zhuo2018message), the particles of SVGD tend to collapse to modes of the target distribution, and this particle degeneracy phenomenon becomes more severe with higher dimensions. This issue was explained in a previous study (ba2021understanding), the authors compared the SVGD update with the gradient descent on the maximum mean discrepancy (MMD). To address this challenge, annealed-SVGD (d2021annealed) employs an annealing schedule to solve the inability of the particles to escape from local modes and the inefficiency in reproducing the density of the different regions. Selecting an appropriate kernel remains a key challenge, as the performance of the algorithm is highly dependent on the choice of kernel (duncan2023geometry). Multiple-kernel learning SVGD (MKL-SVGD) (ai2023stein) enables the combination of multiple kernels into a single, unified kernel, providing greater flexibility in the learning process. In MOO, the first study MOO-SVGD (liu2021profiling) adapted SVGD to profile the Pareto front while maintaining diversity, providing an efficient alternative to traditional methods.
3. Preliminaries
3.1. Multi-objective optimization
@@
(1) $\min_{x\in\mathcal{X}}\mathcal{F}(x)=\begin{bmatrix}f_{1}(x)\\ f_{2}(x)\\ \vdots\\ f_{m}(x)\end{bmatrix}$,
@@
3.2. Pareto Set Learning
Figure 1. The Pareto Set Learning (PSL) framework. (Left) In a general multi-objective problem, the hypernetwork $h(r_{i},\phi)$ maps a preference vector $r_{i}$ to a solution $x_{i}$. (Right) In multi-task learning, the hypernetwork generates the parameters $\theta_{i}$ for a target network $TN(\cdot,\theta_{i})$ that processes data for multiple tasks.
Pareto Set Learning (PSL) offers an efficient approach to obtaining the entire Pareto set in a single training phase, contrasting with traditional methods that require varying hyperparameters to approximate only a few Pareto-optimal points. PSL employs a hypernetwork (ha2016hypernetworks) to generate a diverse set of candidate solutions. During training, the hypernetwork’s parameters are optimized by directly minimizing the multi-objective optimization (MOO) objective. Consequently, the trained hypernetwork produces a dense set of solutions that closely approximate the true Pareto front.
@@ 3.2. Pareto Set Learning
(2) $\begin{split}\phi^{*}=\arg\min_{\phi}\mathbb{E}_{r\sim\text{Dir}(\alpha)}\mathbf{s}\left(\mathcal{F}\left(\mathbf{x}_{r}\right),r\right)\\ \text{s.t. }\mathbf{x}_{r}=\mathbf{h}(r,\phi^{*})\in X_{E},\,\mathbf{h}\left(\mathcal{P},\phi^{*}\right)\equiv X_{E}\end{split}$
[…]
(3) $\phi_{t+1}=\phi_{t}-\xi\nabla_{\phi}\mathbf{s}\left(\mathcal{F}\left(\mathbf{x}_{r}\right),r\right),$
@@
(4) $\phi^{*}=\arg\min_{\phi}\mathbb{E}_{r\sim Dir(\alpha)}s(\mathcal{L}(\cdot,\theta_{r}),r)$
@@
3.3. Stein Variational Gradient Descent (SVGD)
SVGD, a particle-based inference algorithm, is proposed to approximate the target distribution by applying the gradient information in the updating process. SVGD starts at a set of initial points $\{x_{i}^{0}\}_{i=1}^{n}\in\mathbb{R}^{d}$ drawn by a prior distribution $p(x)$, which is pushed to the target points $\{x^{*}_{i}\}_{i=1}^{n}$ belonging to the target distribution $q(x)$, following the iterative:
(5) $\begin{split}x_{i}^{t+1}\leftarrow x_{i}^{t}+\epsilon\phi^{*}(x_{i}^{t}),\forall i=1,...,n,\\ \phi^{*}_{k}=\arg\max_{\phi\ \in\mathcal{B}_{k}}\left\{-\frac{d}{d\epsilon}\mathbf{K}\mathbf{L}(q_{[\epsilon\phi]}||p)\bigg|_{\epsilon=0}\right\}\end{split}$